# Are supernovae driving turbulence in the solar neighborhood?

### 13 Mar 2026 v1

@@ 3.3 Velocity dispersion comparison
After constructing the synthetic sky maps for the different simulations, a velocity cut of $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$ is then applied for a consistent comparison with observations. We then compute the mass-weighted line-of-sight velocity dispersion for each map. Figure 1 shows the obtained distribution of these values. The dot-dashed vertical line corresponds to the median observed H i velocity dispersion, $\sigma_{50}=11.1^{+0.9}_{-0.3}\penalty 10000\ \mathrm{km\,s^{-1}}$, derived in Appendix E from Monte Carlo sampling of LOS velocities reported by S25, using 1000 realizations of the Reid et al. (2019) Galactic rotation model (see Appendix E).
SN-driven simulations produce systematically lower velocity dispersions, with mean values of $\sim 4.9\text{--}6.7\penalty 10000\ \mathrm{km\,s^{-1}}$, indicating that supernova feedback alone underestimates the turbulent level observed in the solar neighborhood. Among these, the SN-1.5 and TIGRESS-MHD4pc runs show the largest spreads, with tails reaching and occasionally exceeding the observational value, suggesting that rare SN-driven configurations can attain the required amplitudes, but that the typical state remains under-turbulent relative to the data.
As for the large–scale–forced runs, they tend to exhibit higher velocity dispersions at larger forcing amplitudes, with values ranging between $\sim 9.6$ and $15.3\penalty 10000\ \mathrm{km\,s^{-1}}$. The velocity dispersion in TURB-20 exceeds the observational reference, which lies between the values measured for TURB-7 and the intermediate forcing runs TURB-10 and TURB-15. The similar dispersions obtained for these two runs, with TURB-10 slightly higher, despite their different forcing amplitudes, are a consequence of the imposed velocity cut, which limits the contribution from the high-velocity tail of the distributions.
@@ 4 Summary and conclusions
We compared solar–neighborhood H i line–of–sight velocities from S25 over the distance range $70\text{--}500$ pc with two families of 1 kpc box simulations that differ in their turbulence drivers: (i) SN–driven runs, where turbulence is powered solely by supernovae, and (ii) large-scale turbulent–forcing runs, where energy is injected at low Fourier modes to mimic the injection from the galactic turbulent cascade.
Our comparison of LOS velocity dispersions shows that the SN-driven simulations systematically underpredict the observed H i dispersion. The typical values obtained in the SN runs fall well below the observational range, $\sigma_{\mathrm{obs}}=10.8\text{--}12\penalty 10000\ \mathrm{km\,s^{-1}}$, inferred from the H i data.
Assuming standard turbulent dissipation, the energy injection rate required to sustain a velocity dispersion $\sigma$ scales as $\dot{E}\propto\sigma^{3}$. This implies that reproducing the observed H i velocity dispersion requires an energy input $\sim 4$$10$ times larger than that provided by SN-driven models, and potentially even more if velocities above the adopted $30\penalty 10000\ \mathrm{km\,s^{-1}}$ $|v_{\mathrm{los}}|$ threshold are taken into account. Additional energy injection on large scales is therefore required. In our setup, we find that a large-scale forcing amplitude between those of TURB-7 and intermediate forcing runs (TURB-10 and TURB-15) is sufficient to reproduce the observed velocity dispersion.
We stress that these estimates should be regarded as lower limits, since the adopted velocity cut is likely conservative and including higher-velocity gas would further increase the measured dispersion. However, identifying the physical nature of the required large-scale driving, will require galactic simulations that include self-consistent galactic-scale physics.
@@ Appendix A Turbulence driving mechanisms
In this appendix, we provide a brief description of the different mechanisms used to drive turbulence in the FRIGG simulations (SN and TURB runs).
@@ A.1 Supernova driving
In this work, the SN runs adopt a time–dependent UV background, and their numerical suffix indicates the initial mid–plane density (in cm-3). Two members of this series, SN-1 and SN-0.66, were presented in Colman et al. (2025) under the names n1_rms00000 and n0.66_rms00000, respectively.
In these runs sink particles (Bate et al. 1995; Federrath et al. 2010b) are introduced when the density exceeds a threshold of $10^{3}\,\mathrm{cm^{-3}}$. These particles can accrete mass from the surrounding gas. Each time sink particles accumulate $120\,\mathrm{M_{\odot}}$ , a stellar particle is created with a random mass between $8$ and $120\,\mathrm{M_{\odot}}$ drawn from a Salpeter initial mass function (Salpeter 1955). The stellar particle is then assigned a main-sequence lifetime related to its mass. Once this lifetime ends, a supernova explosion is triggered, injecting $10^{51}\,\mathrm{erg}$ of thermal energy and $4\times 10^{43}\,\mathrm{g\,cm\,s^{-1}}$ of radial momentum into the surrounding cells (Iffrig and Hennebelle 2015). To avoid very short time steps, a temperature limit ($2\times 10^{6}\,\mathrm{K}$) and a velocity limit ($200\,\mathrm{km\,s^{-1}}$) are set.
@@ A.2 Turbulent forcing
In the TURB runs, turbulence is injected following the Ornstein–Uhlenbeck (OU) process (Eswaran and Pope 1988; Schmidt et al. 2006, 2009; Federrath et al. 2010a) as described in Brucy et al. (2023). At each time step ($dt\simeq 0.4\penalty 10000\ \mathrm{Myr}$), a two–dimensional random acceleration field is evolved in Fourier space, and then transformed back into real space. The forcing is confined to the disc plane and restricted to large–scale Fourier modes with wavelengths between $L_{\rm box}/3$ and $L_{\rm box}$, with the maximum weight assigned to the mode corresponding to a wavelength of $L_{\rm box}/2$. This mimics the injection of kinetic energy from processes at the galactic scale.
We adopt a compressive fraction $\chi=0.5$ (the ratio of compressive to solenoidal modes contributions), which corresponds to the natural mixture of one‑third compressive and two‑thirds solenoidal modes projection (Schmidt et al. 2009; Federrath et al. 2010a). The parameter $f_{\mathrm{rms}}$ controls the power injected by the forcing into the simulation; its value varies in the four runs listed above. Each simulation is evolved for at least two turbulent crossing times, ensuring that the injected turbulence has fully developed throughout the box.
@@ Appendix B Synthetic and observational sky maps
In this appendix, we present a visual comparison between the H i sky map from S25 and an example of the synthetic sky maps generated from the simulations. Figure 2 shows side–by–side the observational map and a full unmasked synthetic map from the TURB-15 run. Figure 3 then shows the four masked realizations obtained by applying the observational sampling pattern after rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ in longitude. These masked maps are the ones used in the velocity–dispersion analysis.
(a) Observed H i line-of-sight velocity map in S25.
(b) Example of one full synthetic line-of-sight velocity map from TURB-15.
Figure 2: Observed and simulated line-of-sight velocity maps before applying the observational mask.
(a) Rotation $0^{\circ}$
(b) Rotation $90^{\circ}$
(c) Rotation $180^{\circ}$
(d) Rotation $270^{\circ}$
Figure 3: Masked synthetic sky maps obtained by applying the observational sampling pattern at different rotation angles to the full synthetic map shown in Fig. 2 (b).
@@ Appendix C Evolution of the velocity PDF
In this appendix, we show how the probability density function (PDF) of the line–of–sight velocity evolves purely as a consequence of applying successive numerical processing steps to the simulations.
@@
  • Panel A: $v_{los}$ PDF of all simulation cells satisfying $70\leq r\leq 500$ pc and $|z|\leq r\sin 5^{\circ}$ in the full 4 pc uniform cube ($\sim 6\times 10^{5}$ cells).
  • Panel B: Same spatial selection, but after collapsing the cube into a 2D $(x,y)$ map by averaging over all cells along the entire vertical line of sight (all $z$ at fixed $x,y$) ($\sim 5\times 10^{4}$ cells).
@@
A clear evolution is visible from Panels A to D: the intrinsic simulation PDF is smooth, but becomes progressively narrower and more irregular as averaging, angular binning, and sky incompleteness are introduced. This shows that features such as a narrow central peak and long high–velocity tails—characteristic of the observational PDF—can arise from numerical sampling effects alone. We do not claim that every synthetic sky reproduces the observed PDF, but the example shown here illustrates how limited resolution and partial coverage can produce these features.
@@
Appendix D Dependencies on the velocity threshold
Figure 5: Evolution of the mass-weighted line-of-sight velocity dispersion as a function of the imposed LOS–velocity threshold for both simulations and observations. Colored curves show the simulated models, averaged over the corresponding sets of synthetic sky maps, with supernova-driven runs shown as colored dotted lines and large-scale turbulent-forcing runs as colored dashed lines. Black curves correspond to the observational H i data from S25: solid for the full dataset, dotted for the outer Galaxy, and dash-dotted for the inner Galaxy.
In Section 3.3, where only velocities up to $30\penalty 10000\ \mathrm{km\,s^{-1}}$ were considered, we showed that SN-driven models yield velocity dispersions below the H i value reported by S25, implying that an additional large-scale driving mechanism is required. The observed values lie between those of the TURB-7 and TURB-10 simulations.
Appendix C further demonstrates that the non-smooth shape of the observed H i line-of-sight velocity PDF, characterized by a sharp central peak and extended high-velocity tails, can be recovered from the simulations once the observational sampling pattern and resolution are applied. This indicates that part of the observed high-velocity structure may arise from resolution and completeness effects, and may reflect the way velocities in these ranges are sampled rather than actual physical anomalies.
@@ Appendix D Dependencies on the velocity threshold
For the observational data, we also separate the inner and outer Galaxy sight lines. Up to $\sim 35\penalty 10000\ \mathrm{km\,s^{-1}}$, both regions present similar behavior, then they diverge after that, with the inner Galaxy showing larger dispersions. At higher velocity thresholds, the continued increase of the observed velocity dispersion is likely dominated by contamination from noise and contributions from distant, non-local high-velocity gas. When comparing simulations with the full Galaxy observational curve, the closest match depends on the chosen threshold: at $30\penalty 10000\ \mathrm{km\,s^{-1}}$ TURB-7 matches best; at $40\penalty 10000\ \mathrm{km\,s^{-1}}$ the closest model becomes TURB-10 or TURB-15; and at $50\penalty 10000\ \mathrm{km\,s^{-1}}$, TURB-20 provides the best agreement. However, reliably determining the forcing amplitude required at these higher thresholds will ultimately require more complete sky coverage and more confident H i measurements at high velocities, in order to distinguish genuine turbulent signal from distant motions or instrumental contamination.
Appendix E Dependencies on the Galactic rotation model
@@
In this appendix, we consider the effects introduced by the uncertainties in the Galactic rotation models employed to subtract Galactic rotation from kinematic tomography line-of-sight velocity estimates in S25. We focus on the consequences of two aspects of the Galactic rotation model. First, we tested the choice of Galactic rotation model by considering alternatives. Second, we evaluated the effect of uncertainties in the parameters of the Galactic rotation model in Reid et al. (2019).
@@
We considered three alternatives to the Reid et al. (2019) model: a flat rotation curve and the models in Brand and Blitz (1993) and Reid et al. (2014). The Reid et al. (2014) model is the previous iteration of the Reid et al. (2019), obtained with 50% fewer parallax measurements. The Brand and Blitz (1993) model is a landmark reference, based on spectro-photometric distances to Hii regions and Hi 21-cm emission observations. Finally, the flat rotation curve, constant circular speed with Galactocentric radius, is an acceptable approximation for the Milky Way’s disk dynamics around the solar radius.
@@
Table 2 presents the standard deviations of the line-of-sight velocities obtained after subtracting the different models of Galactic rotation from the kinematic tomography results in S25. We consider two distance ranges in the input data: 70 <<dd<< 500 pc, which matches the size of the simulation domain, and 70 <<dd<< 1250 pc, which encompasses the distance range in the S25 reconstruction. The results reported in Table 2 demonstrate that there is very little variation in the resulting standard deviation of non-circular motions, as expected from the nearly constant circular speed around the solar radius. The resulting histograms of residual motions, shown in Fig. 6, indicate that the global distributions are very similar for the different Galactic rotation models.
@@
Figure 6: Same as figure 12 in S25 but for three different Galactic rotation models.
@@
Figure 7: Same as figure 12 in S25 but for the sampling of parameters in the (Reid et al. 2019) Galactic rotation model.

### 31 Mar 2026 v2

@@ 3.3 Velocity dispersion comparison
After constructing the synthetic sky maps for the different simulations, a velocity cut of $|v_{\mathrm{los}}|\leq 30\penalty 10000\ \mathrm{km\,s^{-1}}$ was applied for a consistent comparison with observations. We then computed the mass-weighted LOS velocity dispersion for each map. Figure 1 shows the obtained distribution of these values. The dot-dashed vertical line in the figure corresponds to the median observed H i velocity dispersion, $\sigma_{50}=11.1^{+0.9}_{-0.3}\penalty 10000\ \mathrm{km\,s^{-1}}$, derived in Appendix E from Monte Carlo sampling of LOS velocities reported by S25 using 1000 realizations of the Reid et al. (2019) Galactic rotation model (see Appendix E).
The SN-driven simulations produced systematically lower velocity dispersions, with mean values of $\sim 4.9\text{--}6.7\penalty 10000\ \mathrm{km\,s^{-1}}$, indicating that SN feedback alone underestimates the turbulent level observed in the solar neighborhood. Among the SN-driven simulations, the SN-1.5 and TIGRESS-MHD4pc runs show the largest spreads, with tails reaching and occasionally exceeding the observational value, suggesting that rare SN-driven configurations can attain the required amplitudes but that the typical state remains under-turbulent relative to the data.
As for the large-scale–forced runs, they tend to exhibit higher velocity dispersions at larger forcing amplitudes, with values ranging between $\sim 9.6$ and $15.3\penalty 10000\ \mathrm{km\,s^{-1}}$. The velocity dispersion in TURB-20 exceeds the observational reference, which lies between the values measured for TURB-7 and the intermediate forcing runs TURB-10 and TURB-15. The similar dispersions obtained for these two runs (with TURB-10 slightly higher) despite their different forcing amplitudes are a consequence of the imposed velocity cut, which limits the contribution from the high-velocity tail of the distributions.
@@ 4 Summary and conclusions
We have compared solar-neighborhood H i LOS velocities from S25 over the distance range of $70\text{--}500$ pc with two families of 1 kpc box simulations that differ in their turbulence drivers: (i) SN-driven runs, where turbulence is powered solely by SNe, and (ii) large-scale turbulent-forcing runs, where energy is injected at low Fourier modes to mimic the injection from the galactic turbulent cascade. Our comparison of LOS velocity dispersions showed that the SN-driven simulations systematically underpredict the observed H i dispersion. The typical values obtained in the SN runs fall well below the observational range, $\sigma_{\mathrm{obs}}=10.8\text{--}12\penalty 10000\ \mathrm{km\,s^{-1}}$, inferred from the H i data.
Assuming standard turbulent dissipation, the energy injection rate required to sustain a velocity dispersion, $\sigma$, scales as $\dot{E}\propto\sigma^{3}$. This implies that reproducing the observed H i velocity dispersion requires an energy input approximately four to ten times larger than that provided by SN-driven models, and potentially even more if velocities above the adopted $30\penalty 10000\ \mathrm{km\,s^{-1}}$ $|v_{\mathrm{los}}|$ threshold are taken into account. Additional energy injection on large scales is therefore required. In our setup, we find that a large-scale forcing amplitude between those of TURB-7 and intermediate forcing runs (TURB-10 and TURB-15) is sufficient to reproduce the observed velocity dispersion.
We stress that these estimates should be regarded as lower limits since the adopted velocity cut is likely conservative and including higher-velocity gas would further increase the measured dispersion. However, identifying the physical nature of the required large-scale driving will require galactic simulations that include self-consistent galactic-scale physics.
@@ Appendix A Turbulence driving mechanisms
In this appendix we provide a brief description of the different mechanisms used to drive turbulence in the FRIGG simulations. Section A.1 describes how turbulence is driven by SN feedback in the SN runs, while Sect. A.2 explains how turbulence is driven by large-scale forcing in the TURB runs.
@@ A.1 Supernova driving
In this work, the SN runs adopt a time-dependent UV background, and their numerical suffix indicates the initial mid–plane density (in cm-3). Two members of this series, SN-1 and SN-0.66, were presented in Colman et al. (2025) under the names n1_rms00000 and n0.66_rms00000, respectively.
In these runs sink particles (Bate et al. 1995; Federrath et al. 2010b) are introduced when the density exceeds a threshold of $10^{3}\,\mathrm{cm^{-3}}$. These particles can accrete mass from the surrounding gas. Each time sink particles accumulate $120\,\mathrm{M_{\odot}}$ , a stellar particle is created with a random mass between $8$ and $120\,\mathrm{M_{\odot}}$ drawn from a Salpeter initial mass function (Salpeter 1955). The stellar particle is then assigned a main-sequence lifetime related to its mass. Once this lifetime ends, a SN explosion is triggered, injecting $10^{51}\,\mathrm{erg}$ of thermal energy and $4\times 10^{43}\,\mathrm{g\,cm\,s^{-1}}$ of radial momentum into the surrounding cells (Iffrig and Hennebelle 2015). To avoid very short time steps, a temperature limit ($2\times 10^{6}\,\mathrm{K}$) and a velocity limit ($200\,\mathrm{km\,s^{-1}}$) are set.
@@ A.2 Turbulent forcing
In the TURB runs, turbulence is injected following the Ornstein–Uhlenbeck (OU) process (Eswaran and Pope 1988; Schmidt et al. 2006, 2009; Federrath et al. 2010a) as described in Brucy et al. (2023). At each time step ($dt\simeq 0.4\penalty 10000\ \mathrm{Myr}$), a two–dimensional random acceleration field is evolved in Fourier space,and then transformed back into real space. The forcing is confined to the disk plane and restricted to large-scale Fourier modes with wavelengths between $L_{\rm box}/3$ and $L_{\rm box}$, with the maximum weight assigned to the mode corresponding to a wavelength of $L_{\rm box}/2$. This mimics the injection of kinetic energy from processes at the galactic scale.
We adopt a compressive fraction $\chi=0.5$ (the ratio of compressive to solenoidal modes contributions), which corresponds to the natural mixture of one‑third compressive and two‑thirds solenoidal modes projection (Schmidt et al. 2009; Federrath et al. 2010a). The parameter $f_{\mathrm{rms}}$ controls the power injected by the forcing into the simulation; its value varies in the four runs listed above. Each simulation was evolved for at least two turbulent crossing times, ensuring that the injected turbulence has fully developed throughout the box.
@@ Appendix B Synthetic and observational sky maps
In this appendix we present a visual comparison between the H i sky map from S25 and an example of the synthetic sky maps generated from the simulations. Figure 2 shows side–by–side the observational map and a full unmasked synthetic map from the TURB-15 run. Figure 3 then shows the four masked realizations obtained by applying the observational sampling pattern after rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ in longitude. These masked maps are the ones used in the velocity–dispersion analysis.
Figure 2: Observed and simulated LOS velocity maps before applying the observational mask. Panel (a): Observed H i LOS velocity map in S25. Panel (b): Example of one full synthetic LOS velocity map from TURB-15.
Figure 3: Masked synthetic sky maps obtained by applying the observational sampling pattern at different rotation angles to the full synthetic map shown in Fig. 2 (b): Rotation $0^{\circ}$ (panel a), $90^{\circ}$ (b), $180^{\circ}$ (c), and $270^{\circ}$ (d).
@@ Appendix C Evolution of the velocity PDF
In this appendix we show how the probability density function (PDF) of the line–of–sight velocity evolves purely as a consequence of applying successive numerical processing steps to the simulations.
@@
  • Panel a: Point spread function of $v_{los}$ for all simulation cells satisfying $70\leq r\leq 500$ pc and $|z|\leq r\sin 5^{\circ}$ in the full 4 pc uniform cube ($\sim 6\times 10^{5}$ cells).
  • Panel b: Same spatial selection, but after collapsing the cube into a 2D $(x,y)$ map by averaging over all cells along the entire vertical LOS (all $z$ at fixed $x,y$; $\sim 5\times 10^{4}$ cells).
@@
A clear evolution is visible from panels a to d: the intrinsic simulation PDF is smooth, but becomes progressively narrower and more irregular as averaging, angular binning, and sky incompleteness are introduced. This shows that features such as a narrow central peak and long high–velocity tails – characteristic of the observational PDF – can arise from numerical sampling effects alone. We do not claim that every synthetic sky reproduces the observed PDF, but the example shown here illustrates how limited resolution and partial coverage can produce these features.
@@
Appendix D Dependences on the velocity threshold
Figure 5: Evolution of the mass-weighted LOS velocity dispersion as a function of the imposed LOS–velocity threshold for both simulations and observations. Colored curves show the simulated models, averaged over the corresponding sets of synthetic sky maps, with SN-driven runs shown as colored dotted lines and large-scale turbulent-forcing runs as colored dashed lines. Black curves correspond to the observational H i data from S25: solid for the full dataset, dotted for the outer Galaxy, and dash-dotted for the inner Galaxy.
In Sect. 3.3, where only velocities up to $30\penalty 10000\ \mathrm{km\,s^{-1}}$ were considered, we showed that SN-driven models yield velocity dispersions below the H i value reported by S25, implying that an additional large-scale driving mechanism is required. The observed values lie between those of the TURB-7 and TURB-10 simulations.
Appendix C further demonstrates that the non-smooth shape of the observed H i LOS velocity PDF, characterized by a sharp central peak and extended high-velocity tails, can be recovered from the simulations once the observational sampling pattern and resolution are applied. This indicates that part of the observed high-velocity structure may arise from resolution and completeness effects and may reflect the way velocities in these ranges are sampled rather than actual physical anomalies.
@@ Appendix D Dependences on the velocity threshold
For the observational data, we also separate the inner and outer Galaxy sight lines. Up to $\sim 35\penalty 10000\ \mathrm{km\,s^{-1}}$, both regions present similar behavior, then they diverge after that, with the inner Galaxy showing larger dispersions. At higher velocity thresholds, the continued increase of the observed velocity dispersion is likely dominated by contamination from noise and contributions from distant, nonlocal high-velocity gas. When comparing simulations with the full Galaxy observational curve, the closest match depends on the chosen threshold: at $30\penalty 10000\ \mathrm{km\,s^{-1}}$ TURB-7 matches best; at $40\penalty 10000\ \mathrm{km\,s^{-1}}$ the closest model becomes TURB-10 or TURB-15; and at $50\penalty 10000\ \mathrm{km\,s^{-1}}$, TURB-20 provides the best agreement. However, reliably determining the forcing amplitude required at these higher thresholds will ultimately require more complete sky coverage and more confident H i measurements at high velocities, in order to distinguish genuine turbulent signal from distant motions or instrumental contamination.
Appendix E Dependences on the Galactic rotation model
@@
In this appendix we consider the effects introduced by the uncertainties in the Galactic rotation models employed to subtract Galactic rotation from kinematic tomography LOS velocity estimates in S25. We focus on the consequences of two aspects of the Galactic rotation model. First, we tested the choice of Galactic rotation model by considering alternatives. Second, we evaluated the effect of uncertainties in the parameters of the Galactic rotation model in Reid et al. (2019).
@@
We considered three alternatives to the Reid et al. (2019) model: a flat rotation curve and the models in Brand and Blitz (1993) and Reid et al. (2014). The Reid et al. (2014) model is the previous iteration of the Reid et al. (2019), obtained with 50% fewer parallax measurements. The Brand and Blitz (1993) model is a landmark reference, based on spectrophotometric distances to Hii regions and Hi 21-cm emission observations. Finally, the flat rotation curve, constant circular speed with Galactocentric radius, is an acceptable approximation for the Milky Way’s disk dynamics around the solar radius.
@@
Table 2 presents the standard deviations of the LOS velocities obtained after subtracting the different models of Galactic rotation from the kinematic tomography results in S25. We consider two distance ranges in the input data: 70 <<dd<< 500 pc, which matches the size of the simulation domain, and 70 <<dd<< 1250 pc, which encompasses the distance range in the S25 reconstruction. The results reported in Table 2 demonstrate that there is very little variation in the resulting standard deviation of noncircular motions, as expected from the nearly constant circular speed around the solar radius. The resulting histograms of residual motions, shown in Fig. 6, indicate that the global distributions are very similar for the different Galactic rotation models.
@@
Figure 6: Same as Fig. 12 in S25 but for three different Galactic rotation models.
@@
Figure 7: Same as Fig. 12 in S25 but for the sampling of parameters in the (Reid et al. 2019) Galactic rotation model.